7/2x + 1/2x= 3x + 3/2 + 5/2x
The answer they are coming up with is -1, can you show me how the problem is done and how they arrived at that number. Thx

Answer by stanbon(75887)

(7/2)x + (1/2)x = 3x + (3/2) + (5/2)x
---------------------
(8/2)x = (6/2)x + (5/2)x + (3/2)
(8/2)x = (11/2)x + (3/2)
(-3/2)x = 3/2
x = -1
=================
Cheers,
Stan H.